Exercise 13.31 Constants Part A A uniform, solid, 1300.0 kg sphere has a radius of 5.00 m, Find the gravitational force this sphere exerts on a 1.60 kg point mass placed at the following distances from the center of the sphere: (a) 5.05 m, and (b) 2.45 m Submit Request Answer ▼ Part B A2p Submit Request Answer
Answer #1

a)

F = G m1 m2 / r^2

F = 6.67*10^-11* 1300*1.6 / 5.05^2

F = 5.44*10^-9 N

========

b)

since the point is inside the sphereso

F = G m1 m2 r / R^3

F = 6.67*10^-11* 1300*1.6* 2.45 / 5.05^3

F = 1.333*10^-8 N

========